OpenStudy (faiqraees):

Find \[ \int^{1}_{0} \frac{dx}{1+x^3} \] using a reduction formula. @ganeshie8

OpenStudy (eliesaab):

TO check your answer https://www.wolframalpha.com/input/?i=Integrate+1%2F(1%2Bx%5E3)

OpenStudy (eliesaab):

Use \[ \frac{1}{x^3+1}=\frac{2-x}{3 \left(x^2-x+1\right)}+\frac{1}{3 (x+1)} \]

OpenStudy (faiqraees):

I have to make use of a reduction formula